# CMRAG: Co-modality–based visual document retrieval and question answering

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a core paradigm in document question answering tasks. However, existing methods have limitations when dealing with multimodal documents: one category of methods relies on layout analysis and text extraction, which can only utilize explicit text information and struggle to capture images or unstructured content; the other category treats document segmentation as visual input and directly passes it to visual language models (VLMs) for processing, yet it ignores the semantic advantages of text, leading to suboptimal retrieval and generation results. To address these research gaps, we propose Co-Modality–based RAG (CMRAG) framework, which can simultaneously leverage texts and images for more accurate retrieval and generation. Our framework includes two key components: (1) a Unified Encoding Model (UEM) that projects queries, parsed text, and images into a shared embedding space via triplet-based training, and (2) a Unified Co-Modality–informed Retrieval (UCMR) method that statistically normalizes similarity scores to effectively fuse cross-modal signals. To support research in this direction, we further construct and release a large-scale triplet dataset of (query, text, image) examples. Experiments demonstrate that our proposed framework consistently outperforms single-modality–based RAG in multiple visual document question-answering (VDQA) benchmarks. The findings of this paper show that integrating co-modality information into the RAG framework in a unified manner is an effective approach to improving the performance of complex VDQA systems.
@@ 1 Introduction
Large language models (LLMs) have received extensive attention in recent years (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025; Yang et al., 2025a), but they have inherent limitations in handling out-of-domain knowledge (Ji et al., 2023). To address this issue, RAG integrates external knowledge retrieval with the generation process (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020; Chen et al., 2025). RAG achieves wide success in open-domain question answering, knowledge retrieval, and dialogue systems, and becomes an effective means of extending the knowledge boundaries of LLMs (Ram et al., 2023; Gao et al., 2023). However, most external data sources (e.g., documents) are essentially multimodal (Jeong et al., 2025; Faysse et al., 2025), often containing natural language text, formulas, tables, images, and complex layout structures. How to effectively leverage such multimodal information in question answering remains a challenging problem that is not fully solved.
One line of approaches is text-based RAG, which typically relies on layout parsing and text extraction (Xu et al., 2020; Dong et al., 2025; Yang et al., 2025b; Perez & Vizcaino, 2024). These methods first detect document layouts and then extract textual information for subsequent retrieval and generation, as shown in Fig. 1(a). While stable at the semantic level, they struggle to handle content such as images and tables. Recently, VLMs (Ghosh et al., 2024; Radford et al., 2021; Alayrac et al., 2022) enable RAG systems to process documents directly as images (Faysse et al., 2025; Yu et al., 2025; Qi et al., 2024a; Wang et al., 2025b), giving rise to vision-based RAG (Huang et al., 2022; Kim et al., 2022; Yu et al., 2025). Specifically, as shown in Fig. 1(b), these methods divide document pages into image segments and perform retrieval and reasoning through visual understanding models. Although they capture non-textual information, they often overlook the precise information carried by text, leading to performance bottlenecks.
@@ 2 Related work
In this section, we discuss the recent studies related to multi-modal RAG (MMRAG). Specifically, we primarily focus on three folds: (1) knowledge–based MMRAG, (2) video–based MMRAG, and (3) document–based MMRAG.
Knowledge–based MMRAG refers to retrieving knowledge (text or image modality) from external sources such as Wikipedia articles and websites to answer textual or visual queries (Talmor et al., 2021; Marino et al., 2019; Chang et al., 2022; Schwenk et al., 2022; Mensink et al., 2023; Chen et al., 2023; Ma et al., 2024a; Hu et al., 2025). Although the external knowledge database can enhance the system performance (Caffagni et al., 2024), the key issue of knowledge–based MMRAG is the inconsistency between textual and visual queries as well as the external knowledge database (Chen et al., 2022; Lin et al., 2023; Zhang et al., 2024b). To address this issue, Lin & Byrne (2022) adopted multiple algorithms, including object detection, image captioning, and OCR, to transform visual queries into language space, and proposed a joint training scheme to optimize retrieval and generation simultaneously. A similar training strategy was also used by (Adjali et al., 2024). Also, Yan & Xie (2024) used a consistent modality for both retrieval and generation: visual modality for retrieval (visual queries and Wikipedia article pages) and textual modality for generation (textual queries and wiki articles). A similar strategy can also be found in RORA (Qi et al., 2024a). In addition, Tian et al. (2025) proposed cross-source knowledge reconciliation for MMRAG, which could address the inconsistency between textual and visual external knowledge.
Video–based MMRAG refers to retrieving videos from the corpus to help answer given queries (Caba Heilbron et al., 2015; Xu et al., 2016; Anne Hendricks et al., 2017; Wang et al., 2019; Kriz et al., 2025; Wan et al., 2025). Since encoding videos may incur high computational costs, a few studies pre-processed videos using VLMs and converted videos to textual modality (Zhang et al., 2024a; Arefeen et al., 2024; Ma et al., 2025). For example, Zhang et al. (2024a) first detected key information in videos such as human faces, based on which a VLM was prompted to generate scene captions for video frames. Consequently, the video modality can be converted to a text modality, which can significantly reduce computational costs and facilitate retrieval and generation. Furthermore, a few studies (Luo et al., 2024; Jeong et al., 2025; Reddy et al., 2025) processed videos by selecting or clustering representative frames, facilitating video retrieval and final generation.
Document–based MMRAG refers to retrieving a few document pages from one or multiple documents to help generate answers for given questions (Methani et al., 2020; Masry et al., 2022; Tanaka et al., 2023; Tito et al., 2023; Ma et al., 2024b; Li et al., 2024; Hui et al., 2024; Qi et al., 2024b; Cho et al., 2024; Wang et al., 2025a; Li et al., 2025; Wasserman et al., 2025; Faysse et al., 2025). Traditionally, documents were parsed using detection models (Ge et al., 2021) and OCR engines (Smith, 2007), and the extracted components (e.g., text) were input to LLMs to generate answers (Riedler & Langer, 2024; Faysse et al., 2025). With the proliferation of VLMs, a few studies (Faysse et al., 2025; Yu et al., 2025; Wang et al., 2025b) processed document pages as images directly. Specifically, they used VLMs to encode queries and document pages as text and images, respectively, based on which the similarity scores between queries and visual document pages can be calculated. This method paves the way for document–based MMRAG, as it does not need to parse documents and can retrieve document pages directly. However, this method overlooks text modality in documents, which may degrade the performance of RAG systems.
@@ 3 Methodology
In this section, we introduce the problem definition, the proposed CMRAG framework, the design and training method of the CMRAG-retrieval, and computational cost analysis.
@@ 3.1 Problem definition
We formulate the task of Visual Document Question Answering (VDQA) within a Retrieval–Augmented Generation (RAG) framework. A collection of visual documents (e.g., PDFs, scanned articles) serves as the knowledge source for answering user queries. As illustrated in Fig. 2(a), each document page $p_{i}$ is first parsed by a Vision Language Model (VLM) $\mathcal{V}$ to extract its structured multimodal content. This parsing step produces both a visual representation $I_{i}$ (which represents the entire page image in this study) and a textual representation $T_{i}$, such that ${I_{i},T_{i}}=\mathcal{V}(p_{i})$. The complete set of all parsed pages constitutes the candidate evidence pool, denoted as $\mathcal{D}=\{\{I_{1},T_{1}\},\{I_{2},T_{2}\},\dots,\{I_{M},T_{M}\}\}$, where $M$ is the total number of pages. For a given query $q$, the objective of the retriever $\mathcal{R}$ is to identify the top-$k$ most relevant pages $P_{k}$ based on a multimodal similarity score $s(q,\{I_{i},T_{i}\})$.
Considering recent advances in VLMs (Bai et al., 2025; Du et al., 2025) enable them to directly process multiple naive-resolution page images, the retrieved evidence $P_{k}$ is directly combined with the query $q$ into a structured prompt $\mathcal{P}(q,P_{k})$, which is then fed into a generator model $\mathcal{G}$, typically a powerful VLM, to produce the final answer: $\hat{a}=\mathcal{G}(\mathcal{P}(q,P_{k}))$ (as shown in Fig. 2(d)). A comprehensive list of notations is provided in Appendix A. Also, we introduce all the prompt templates used in this study in Appendix B.
@@ 4 Data Construction
Training data. Our training corpus is built upon the synthetic dataset from VisRAG (Yu et al., 2025), which comprises diverse PDF documents (e.g., textbooks, academic papers, manuals) and approximately 240k query-document pairs generated by GPT-4o. To fully exploit the multimodal nature of these documents, we re-process all document pages (around 40k) using Qwen2.5-VL-7B (Bai et al., 2025) for end-to-end document parsing. This step extracts a structured representation for each page, including the entire page image, segmented sub-figures, and OCR-based text stored in HTML format. This enriched dataset provides fine-grained, aligned multimodal supervision, which is crucial for training our model to perform effective cross-modal retrieval and reasoning. Detailed data sources and processing prompts are provided in Appendix E.
Evaluation data. To comprehensively evaluate our method, we adopt several widely used VDQA benchmarks that span diverse domains and task types. Specifically, we include MMLongBench (Ma et al., 2024b) [^1] https://huggingface.co/datasets/MMlongBench/MMLongBench-Doc, LongDocURL (Deng et al., 2024) [^2] https://huggingface.co/datasets/THUDM/LongDocURL, and REAL-MM-RAG (Wasserman et al., 2025) [^3] https://huggingface.co/collections/ibm-research. These datasets cover industrial documents, scientific papers, presentation slides, and long multi-page documents, providing a broad and challenging testbed for assessing the retrieval and generation capabilities of our model. Detailed statistics of the evaluation datasets are provided in Tab. 1.
@@ 5.1 Experiment setup
Baselines. We evaluate the effectiveness of our proposed method against several strong embedding model baselines: (1) BGE: A state-of-the-art text embedding model that uses only the parsed textual content from the documents (Xiao et al., 2023); (2) CLIP-B/32: The CLIP model with a Vision Transformer-B/32 image encoder; (3) CLIP-L/14-336: A larger and higher-resolution CLIP variant with a Vision Transformer-L/14 image encoder pretrained on $\times336$ pixel images (Radford et al., 2021); (4) SigLIP: The SigLIP model, which forms the foundation of our architecture (Zhai et al., 2023); and (5) SigLIP2: The latest iteration of the SigLIP model (Tschannen et al., 2025).
@@ 5.4 Analysis
Case study. As illustrated in Fig. 4, a case study from finreport further demonstrates the practical effectiveness of our CMRAG framework. For a given query, the ground truth (GT) evidence is located on page 33. Our CMRAG retriever correctly identifies this GT page as the most similar result. On the contrary, the SigLIP baseline, which relies solely on the image modality, ranks the GT page only as the second most similar. Furthermore, even when the GT page is provided directly to the generator, the baseline fails to produce an accurate answer, whereas our method succeeds. This suggests that for pages with dense textual content, treating them as pure images is insufficient for accurate comprehension. The generator requires the explicit textual information that our method provides. This case underscores the dual advantage of CMRAG: it improves both retrieval accuracy and the quality of the final generated answer. We provide more cases in Appendix G.
Figure 4: A case comes from Finreport.

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a core paradigm in document question answering tasks. However, existing methods have limitations when dealing with multimodal documents: one category of methods relies on layout analysis and text extraction, which can only utilize explicit text information and struggle to capture images or unstructured content; the other category treats document segmentation as visual input and directly passes it to visual language models (VLMs) for processing, yet it ignores the semantic advantages of text, leading to suboptimal retrieval and generation results. To address these research gaps, we propose Co-Modality–based RAG (CMRAG) framework, which can simultaneously leverage texts and images for more accurate retrieval and generation. Our framework includes two key components: (1) a Unified Encoding Model (UEM) that projects queries, parsed text, and images into a shared embedding space via triplet-based training, and (2) a Unified Co-Modality–informed Retrieval (UCMR) method that statistically normalizes similarity scores to effectively fuse cross-modal signals. To support research in this direction, we further construct and release a large-scale triplet dataset of (query, text, image) examples. Experiments demonstrate that our proposed framework consistently outperforms single-modality–based RAG in multiple visual document question-answering (VDQA) benchmarks. The findings of this paper show that integrating co-modality information into the RAG framework in a unified manner is an effective approach to improving the performance of complex VDQA systems. The code is available at https://github.com/ChenWangHKU/CMRAG.
@@ 1 Introduction
Large language models (LLMs) have received extensive attention in recent years (Touvron et al., 2023; Achiam et al., 2023; Guo et al., 2025; Yang et al., 2025a), but they have inherent limitations in handling out-of-domain knowledge (Ji et al., 2023). To address this issue, RAG integrates external knowledge retrieval with the generation process (Lewis et al., 2020; Guu et al., 2020; Karpukhin et al., 2020; Chen et al., 2025; 2026). RAG achieves wide success in open-domain question answering, knowledge retrieval, and dialogue systems, and becomes an effective means of extending the knowledge boundaries of LLMs (Ram et al., 2023; Gao et al., 2023). However, most external data sources (e.g., documents) are essentially multimodal (Jeong et al., 2025; Faysse et al., 2025; Yu et al., 2025b), often containing natural language text, formulas, tables, images, and complex layout structures. How to effectively leverage such multimodal information in question answering remains a challenging problem that is not fully solved.
One line of approaches is text-based RAG, which typically relies on layout parsing and text extraction (Xu et al., 2020; Dong et al., 2025; Yang et al., 2025b; Perez and Vizcaino, 2024). These methods first detect document layouts and then extract textual information for subsequent retrieval and generation, as shown in Fig. 1(a). While stable at the semantic level, they struggle to handle content such as images and tables. Recently, VLMs (Ghosh et al., 2024; Radford et al., 2021; Alayrac et al., 2022) enable RAG systems to process documents directly as images (Faysse et al., 2025; Yu et al., 2025a; Qi et al., 2024a; Wang et al., 2025b), giving rise to vision-based RAG (Huang et al., 2022; Kim et al., 2022; Yu et al., 2025a). Specifically, as shown in Fig. 1(b), these methods divide document pages into image segments and perform retrieval and reasoning through visual understanding models. Although they capture non-textual information, they often overlook the precise information carried by text, leading to performance bottlenecks.
@@ 2 Related work
Document–based MMRAG refers to retrieving a few document pages from one or multiple documents to help generate answers for given questions (Methani et al., 2020; Masry et al., 2022; Tanaka et al., 2023; Tito et al., 2023; Ma et al., 2024b; Li et al., 2024; Hui et al., 2024; Qi et al., 2024b; Cho et al., 2024; Wang et al., 2025a; Li et al., 2025; Wasserman et al., 2025; Faysse et al., 2025). Traditionally, documents were parsed using detection models (Ge et al., 2021) and OCR engines (Smith, 2007), and the extracted components (e.g., text) were input to LLMs to generate answers (Riedler and Langer, 2024; Faysse et al., 2025). With the proliferation of VLMs, a few studies (Faysse et al., 2025; Yu et al., 2025a; Wang et al., 2025b) processed document pages as images directly. Specifically, they used VLMs to encode queries and document pages as text and images, respectively, based on which the similarity scores between queries and visual document pages can be calculated. This method paves the way for document–based MMRAG, as it does not need to parse documents and can retrieve document pages directly. However, this method overlooks text modality in documents, which may degrade the performance of RAG systems. We also discuss knowledge–-base and video–based MMRAG papers in Appendix C.
@@ 3.1 Problem definition
We formulate the task of Visual Document Question Answering (VDQA) within a Retrieval–Augmented Generation (RAG) framework. A collection of visual documents (e.g., PDFs, scanned articles) serves as the knowledge source for answering user queries. As illustrated in Fig. 2(a), each document page $p_{i}$ is first parsed by a Vision Language Model (VLM) $\mathcal{V}$ to extract its structured multimodal content. This parsing step produces both a visual representation $I_{i}$ (which represents the entire page image in this study) and a textual representation $T_{i}$ (which refers to the extracted text from the page), such that ${I_{i},T_{i}}=\mathcal{V}(p_{i})$. The complete set of all parsed pages constitutes the candidate evidence pool, denoted as $\mathcal{D}=\{\{I_{1},T_{1}\},\{I_{2},T_{2}\},\dots,\{I_{M},T_{M}\}\}$, where $M$ is the total number of pages. For a given query $q$, the objective of the retriever $\mathcal{R}$ is to identify the top-$k$ most relevant pages $P_{k}$ based on a multimodal similarity score $s(q,\{I_{i},T_{i}\})$. In this paper, we use the entire page image
Considering recent advances in VLMs (Bai et al., 2025; Du et al., 2025) enable them to process multiple native-resolution page images directly, the retrieved evidence $P_{k}$ is directly combined with the query $q$ into a structured prompt $\mathcal{P}(q,P_{k})$, which is then fed into a generator model $\mathcal{G}$, typically a powerful VLM, to produce the final answer: $\hat{a}=\mathcal{G}(\mathcal{P}(q,P_{k}))$ (as shown in Fig. 2(d)). A comprehensive list of notations is provided in Appendix A. Also, we introduce all the prompt templates used in this study in Appendix B.
@@ 4 Data Construction
Training data. Our training corpus is built upon the synthetic dataset from VisRAG (Yu et al., 2025a), which comprises diverse PDF documents (e.g., textbooks, academic papers, manuals) and approximately 240k query-document pairs generated by GPT-4o. To fully exploit the multimodal nature of these documents, we re-process all document pages (around 40k) using Qwen2.5-VL-7B (Bai et al., 2025) for end-to-end document parsing. This step extracts a structured representation for each page, including the entire page image, segmented sub-figures, and OCR-based text stored in HTML format. This enriched dataset provides fine-grained, aligned multimodal supervision, which is crucial for training our model to perform effective cross-modal retrieval and reasoning. Detailed data sources and processing prompts are provided in Appendix E.
Evaluation data. To comprehensively evaluate our method, we adopt several widely used VDQA benchmarks that span diverse domains and task types. Specifically, we include MMLongBench (Ma et al., 2024b), LongDocURL(Deng et al., 2024), and REAL-MM-RAG (Wasserman et al., 2025). These datasets cover industrial documents, scientific papers, presentation slides, and long multi-page documents, providing a broad and challenging testbed for assessing the retrieval and generation capabilities of our model. Detailed statistics of the evaluation datasets are provided in Tab. 1.
@@ 5.1 Experiment setup
Baselines. We evaluate the effectiveness of our proposed method against several strong embedding model baselines: (1) BGE: A state-of-the-art text embedding model that uses only the parsed textual content from the documents (Xiao et al., 2023); (2) CLIP-B/32: The CLIP model with a Vision Transformer-B/32 image encoder; (3) CLIP-L/14-336: A larger and higher-resolution CLIP variant with a Vision Transformer-L/14 image encoder pretrained on $\times336$ pixel images (Radford et al., 2021); (4) SigLIP: The SigLIP model, which forms the foundation of our architecture (Zhai et al., 2023); and (5) SigLIP2: The latest iteration of the SigLIP model (Tschannen et al., 2025). It should be noted that this study implements and trains the embedding model based on SigLIP, rather than advanced VLMs. Hence, we focus on comparing the performance between UEM and the existing embedding models with a comparable model size, while excluding those models with a dramatically heavy architecture (Yu et al., 2025a; Faysse et al., 2025; Wang et al., 2025b).
@@ 6 Conclusion
Beyond VDQA benchmarks, CMRAG is applicable to a broad range of document retrieval–centric settings where evidence is inherently multimodal and neither OCR-only text nor pure vision chunks are sufficient. Typical use cases include (i) enterprise knowledge search over slide decks, reports, manuals, and scanned PDFs, where users ask for procedures, specifications, or comparisons that may be expressed in surrounding text but grounded by figures, tables, or screenshots; (ii) technical troubleshooting and customer support, where queries reference UI screenshots, wiring diagrams, error plots, or configuration tables and effective retrieval must align textual descriptions with visual cues; and (iii) scientific and educational document assistance, where answers depend on jointly interpreting captions, formulas, plots, and explanatory paragraphs. In all these scenarios, the key challenge is cross-modal evidence matching and fusion under heterogeneous similarity-score distributions—precisely what UEM’s unified embedding space and UCMR’s normalized score fusion are designed to address—making CMRAG a general retrieval-and-grounded-generation backbone for practical multimodal document applications.
@@ Appendix C Additional related work
Knowledge–based MMRAG refers to retrieving knowledge (text or image modality) from external sources such as Wikipedia articles and websites to answer textual or visual queries (Talmor et al., 2021; Marino et al., 2019; Chang et al., 2022; Schwenk et al., 2022; Mensink et al., 2023; Chen et al., 2023; Ma et al., 2024a; Hu et al., 2025). Although the external knowledge database can enhance the system performance (Caffagni et al., 2024), the key issue of knowledge–based MMRAG is the inconsistency between textual and visual queries as well as the external knowledge database (Chen et al., 2022; Lin et al., 2023; Zhang et al., 2024b). To address this issue, (Lin and Byrne, 2022) adopted multiple algorithms, including object detection, image captioning, and OCR, to transform visual queries into language space, and proposed a joint training scheme to optimize retrieval and generation simultaneously. A similar training strategy was also used by (Adjali et al., 2024). Also, (Yan and Xie, 2024) used a consistent modality for both retrieval and generation: visual modality for retrieval (visual queries and Wikipedia article pages) and textual modality for generation (textual queries and wiki articles). A similar strategy can also be found in RORA (Qi et al., 2024a). In addition, (Tian et al., 2025) proposed cross-source knowledge reconciliation for MMRAG, which could address the inconsistency between textual and visual external knowledge.
@@ Appendix F Retrieval experiment results
Appendix G Case study
As illustrated in Fig. 6, a case study from finreport further demonstrates the practical effectiveness of our CMRAG framework. For a given query, the ground truth (GT) evidence is located on page 33. Our CMRAG retriever correctly identifies this GT page as the most similar result. On the contrary, the SigLIP baseline, which relies solely on the image modality, ranks the GT page only as the second most similar. Furthermore, even when the GT page is provided directly to the generator, the baseline fails to produce an accurate answer, whereas our method succeeds. This suggests that for pages with dense textual content, treating them as pure images is insufficient for accurate comprehension. The generator requires the explicit textual information that our method provides. This case underscores the dual advantage of CMRAG: it improves both retrieval accuracy and the quality of the final generated answer.
Figure 6: A case comes from Finreport.